40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 10 19:03:57 2011, rana, Update, LSC, Y Arm work 
    Reply  Sat Sep 10 19:45:29 2011, Jenne, Update, LSC, Y Arm work 
    Reply  Tue Sep 13 16:57:44 2011, steve, Update, LSC, ITMX oplev ** 
Message ID: 5381     Entry time: Sat Sep 10 19:03:57 2011     Reply to this: 5382   5395
Author: rana 
Type: Update 
Category: LSC 
Subject: Y Arm work 

I lined up the Y Arm for locking and then centered the oplevs for ETMY and ITMY.

* The ITMY OL has still got the old style laser. Steve, pleaes swap this one for a HeNe. Also the optical layout seems strange: there are two copies of the laser beam going into the chamber (??). Also, the QPD transimpedance needs to be increase by a factor of ~10. We're only getting ~500 counts per quadrant. Its worth it for someone to re-examine the whole ITMY OL beam layout.

* The ETMY OL beam was coming out but clipping on the mount for the ETMY OL HeNe. This indicates a failure on our part to do the ETMY closeout alignment properly. In fact, I get the feeling from looking around that we overlooked aligning the OL and IPPOS/ANG beams this time. If we're unlucky this could cause us to vent again. I undid part of the laser mount and changed the height on the receiving mirror to get the beam back onto the QPD.

I noticed that there is significant green light now getting into some of the IR PDs; beacuse of this there are weird offsets in the TRY QPD and perhaps elsewhere. We had better purchase some filters to tape over the front of the sensitive IR sensors to prevent the couplling from the green laser.

* There is a beam on IPPOS, but its too big for the detector (this has always been the case). We need to put a 2" lens with a weak focusing power on this path so as to halve the beam size on the detector. Right now its clipping and misleading. There is also a 0.9V offset on the SUM signal. I'm not sure if this readout is working at all.

* I couldn't find any beam on IPANG at all. Not sure what's changed since Kiwamu saw it.

ELOG V3.1.3-